OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x – y – 3z = –3 x – 2y – 2z = –3 x + y + 3z = –2 (–1, 2, –1) (1, –2, 1) (–1, –2, 1) (1, –2, –1)

OpenStudy (anonymous):

ok subtract equation 3 from 2 x-2y-2z-(x+y+3z)=-3-(-2) -3y-5z=-1 ...........A multiply the equation 2 by 2 4x-4y-8z=-6 subtract this from equation 1 4x-y-3z-(4x-4y-8z)=-3-(-6) 3y+5y=3 ..............B solve equations A and B

OpenStudy (anonymous):

@priscillaam_0422 can you solve A and B equations ?

OpenStudy (anonymous):

yes! thank you somuch!

OpenStudy (anonymous):

and there is a typo the Equation B is 3y+5z=3 ..............B

OpenStudy (anonymous):

yw:)